OpenStudy (anonymous):

Ms.Lewis plans to drive 900 miles to her vacation destination, driving an average of 50 miles per hour. How many miles per hour faster must she average, while driving, to reduce her total driving time by 3 hours?

OpenStudy (anonymous):

If she were to go 50 mph, the trip would have taken her 900/50 or 18 hours. She wishes, then, to go 15 hours instead, so 900/x = 15 And just solve for x.

OpenStudy (anonymous):

The mph "faster" will be x - 50

OpenStudy (anonymous):

r t = d Solve the following for x, the additional speed required to reduce the time by 3 hours.\[(50+x)\left(\frac{900}{50}-3\right)=900\]x = 10 mph